OpenStudy (anonymous):

write the equation of the circle whose center is at (4,-9) and passes through (-2,5)

OpenStudy (triciaal):

equation of a circle (x-x1)^2 + (y - y2)^2 = r^2 where the center is (x1,y1) and (x, y) lies on the circle

OpenStudy (anonymous):

you lost me after equation of circle

OpenStudy (triciaal):

 when the center of the circle is (0, 0) the equation is x^2 + y^2 = r^2 the x coordinate is approximately the radius r and pythagorean theorem